Circular No. 5724

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

X OPHIUCHI A. M. Le Squeren and S. Etoka, Observatoire de Paris, Meudon, report:

“In the course of an OH-maser-emission monitoring of Mira stars with the Nançay Radio Telescope we have observed a strong flare at 1665 MHz in the binary Mira star X Oph. From 1985 to 1991 only weak and fleeting 1667-MHz emission was present near –55.5 km/s (with respect to the local standard of rest) with a flux of < 0.7 Jy. In early 1992 we suddenly observed increasing 1665-MHz polarized emission centered at about –56.7 km/s. The peak flux density we recorded in May 1992 reached 6.8 and 1.8 Jy for righthanded and left-handed circular polarization, respectively. The velocity of this flare emission is close to the –57.4 km/s of the H_{2}O maser. The peak flux density has since decreased, but it is still at the level of 3.5 (righthanded) and 1.2 Jy (left-handed). Optical, OH radio interferometry and other maser-line observations would be useful in order to understand the peculiar nature of this outburst. A similar phenomenon was observed by Le Squeren and Sivagnanam in R Leo in 1985 (IAUC 4164).”

1992 WD_{5}

Improved orbital elements from MPC 21977:

Epoch = 1993 Aug. 1.0 TT

T = 1993 Feb. 13.8403 TT  ω = 122.^{o}1982
e = 0.303944  Ω = 10.1944 }2000.0
q = 1.248736 AU  i = 6.8747
a = 1.794017 AU n^{o} = 0.4101703 P = 2.403 years
1993 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  V
Mar. 24 8^{h}06.^{m}85 +28^{o}34.^{′}3 0.506 1.289 114.^{o}1 44.^{o}9 15.8
Apr. 3 8 34.90 +24 26.7 0.559 1.312 111.5 45.2 16.1
13 9 02.13 +20 23.4 0.621 1.339 108.9 45.1 16.3
23 9 28.37 +16 28.2 0.690 1.370 106.2 44.8 16.6
May 3 9 53.70 +12 43.0 0.768 1.403 103.5 44.3 16.9
13 10 18.13 + 9 08.8 0.854 1.439 100.7 43.6 17.1
23 10 41.79 + 5 45.8 0.946 1.477  97.8 42.8 17.4
June 2 11 04.83 + 2 33.3 1.046 1.516  94.7 41.8 17.6
12 11 27.34 − 0 29.1 1.152 1.556  91.6 40.7 17.9
22 11 49.45 − 3 22.0 1.263 1.597  88.2 39.5 18.1

1993 March 25 Brian G. Marsden